OpenStudy (anonymous):

use implicit integration to find dy/dx . y sin(1/y)= 1-xy

OpenStudy (inkyvoyd):

use the chain rule, and treat y as a differentiabe function of x.

OpenStudy (amistre64):

i see chain rule, product rule, trig stuffs, ... yep

OpenStudy (amistre64):

implicit is no different from explicit; derivative rules are the same regardless of what they want to name it

OpenStudy (anonymous):

can you show me please?

OpenStudy (amistre64):

show you the rules of derivatives?

OpenStudy (anonymous):

no. how to solve it.

OpenStudy (amistre64):

i spose i could ... what the derivative of: xy ?

OpenStudy (anonymous):

y i think. it said we must search dy/ dy by using implicit integration

OpenStudy (amistre64):

its the product rule, can you show me the generic derivative of a product?

OpenStudy (anonymous):

what is that. i'm sorry but i don;t know

OpenStudy (amistre64):

if you dont know what the product rule is for derivatives, then you are asking the wrong question ....

OpenStudy (anonymous):

i know what is product rule.

OpenStudy (amistre64):

(fg)' = f'g+fg'

OpenStudy (amistre64):

notations differ which is why i want you to write up what you know it as

OpenStudy (anonymous):

\[y \sin (1/y) = 1-xy\] find thedy/dx by using the implicit derivatives.

OpenStudy (amistre64):

yes, thats what you wrote to begin with and at the moment we need to work it out, piece by piece so that you get a better concept of it

OpenStudy (amistre64):

lets start with the simplest part of it and derive: xy

OpenStudy (anonymous):

x dy/dx+ y. right??

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

and the derivative of 1 is always 0 so that takes care of the right side. y sin(1/y) = xy' + y the left side is a product rule as well

OpenStudy (anonymous):

i know the concept already. but to derive ysin (1/y) i stuck already

OpenStudy (amistre64):

forget the we are looking at a sin(...) and product it to begin with

OpenStudy (amistre64):

[y sin(1/y)]' = y' sin(1/y) + y sin'(1/y) thats simple enough; now we can focus on what sin'(1/y) needs to be

OpenStudy (amistre64):

you agree with that so far?

OpenStudy (anonymous):

yes yes. but then when to put dy/dx afer derive the 1/y?

OpenStudy (amistre64):

when? the same time you do it normally :) so far we have this \[\frac{dy}{dx}sin(\frac{1}{y})+y\frac{d}{dx}(sin(\frac{1}{y}))=x\frac{dy}{dx}+y\] the final is a chain rule \[\frac{d}{dx}(sin(\frac{1}{y}))\to\ cos(\frac{1}{y})\ \frac{d}{dx}(\frac{1}{y})\] \[\frac{d}{dx}(\frac{1}{y})\to\ \frac{dy}{dx}(\frac{-1}{y^2})\]

OpenStudy (amistre64):

apparently, someone thought it a good idea to have the switch for the wireless internet to be placed on the keypad itself so that when you hit it by accident; you get disconnected :/

OpenStudy (anonymous):

what do you mean?

OpenStudy (amistre64):

i mean what i said. my wireless internet switch is in a bad spot.

OpenStudy (anonymous):

mine too. never mind.

OpenStudy (amistre64):

:) all those dy/dx stuff crowd the scene, i like the newton notation better for this: \[y'\ sin(\frac{1}{y}) + -(\frac{y}{y^2}) cos(\frac{1}{y})\ y' = xy' + y\] the rest is just algebra to isolate the y' s

OpenStudy (anonymous):

ok. got it. thank you. :)

OpenStudy (amistre64):

youre welcome; and good luck